OpenStudy (anonymous):

help me pleasee! The slope of a line is –4 and its y-intercept is (0, 4). What is the equation of the line that is perpendicular to the first line and passes through (–8, 2)?

OpenStudy (sheraz12345):

m=-4 and c=4.... Nnow line is perpendicular to first one so its slope will be 1/4...

OpenStudy (anonymous):

(y=1/4x+4) or (y=1/4x-4)

OpenStudy (anonymous):

@Sheraz12345

OpenStudy (sheraz12345):

No mate u got me wrong .. I only told u the slope.. Now apply formula. y – y1 = m(x – x1)